# VR-GS: A Physical Dynamics-Aware Interactive Gaussian Splatting System in Virtual Reality

### v1

@@ 7. User Study
We conducted a user study with 10 participants: 2 professionals (P1–P2) with 5 and 7 years of experience in 3D VFX software (including Houdini and Blender), and 8 novices (P3–P10), with P2 also having 2 years of VR experience. Our study used the hardware specified in the experimental section and included the Fox, Ring Toss, and Toy Collection demos. Participants received a 10-minute video tutorial on our system before completing two tasks. Post-task, they answered a questionnaire covering usability (System Usability Scale (Bangor et al., 2009)) and subjective feedback on individual (M1-M5: object manipulation, inpainted scene, illumination, realistic dynamics, physics-based placement) and overall system features (M6-M9: easiness to use, latency satisfaction, functionality, overall satisfaction), as shown in Section 7.2. Additionally, we conducted a 20-minute semi-structured interview for more in-depth feedback.
@@ 7.1. Tasks
The participants were required to pet a fox (Figure 7) and complete two VR games developed from our system, including ring toss (Figure 9) and toy collection (Figure 10). Each subtask is a 10-minute session. More specifically, in the ring toss task, users are required to throw rings over the objects on the table, and in the toy collection task, users are required to collect messy toys from the living room into a basket. Task 1 has been specifically designed to enable users to interact comprehensively with the primary features of our system. The order of the three subtasks was counterbalanced by a Latin square design to minimize possible learning effects.
@@
The participants are asked to edit the scene and generate dynamics freely with our prototype system, which includes geometry-based editing, physics-based editing, transform, rotation, duplicating, undoing, rescaling, etc.
@@ 7.2. Results and Discussion
The inset figure summarizes the subjective feedback about the individual features and system features including easiness to use, latency satisfaction, system functionality, and overall satisfaction. Overall, our system has received positive opinions. Among various features, users spoke highly of the placement of virtual content. Transforming objects without physics rules resulted in floating objects, while in our system, all virtual content placements, such as putting toys onto a sofa, are followed by physics rules and are consistent with the user’s understanding and convention. P7 commented, ”I love this a lot. I have a dog. Petting the fox is really like what I did to my dog at home. I felt so real. Moreover, the placement of the toy is awesome. After they all fell to the basket, the basket even shook for a while” The VR-GS system’s realistic lighting and physics-based dynamics create an immersive experience. As a professional user, P2 spoke highly of high-fidelity generative dynamics and illumination. ”Those motions of virtual content are just like what I see in the physical world. I can’t wait to take photos of my own house and then put them in VR with my video game character.” Users rated the system highly on ease of use (4.6/5) and overall satisfaction (4.8/5). With a System Usability Scale (SUS) score of 83.5, VR-GS is classified as ’excellent’ according to (Bangor et al., 2009).
@@ 8. Conclusion and Future Work
This paper has presented VR-GS, a physical dynamics-aware interactive Gaussian Splatting system in VR, particularly first addressing the challenge of editing real-time high-fidelity virtual content. By leveraging the advancements in Gaussian Splatting, VR-GS bridges the quality gap traditionally observed between machine-generated and manually created 3D content. Our system not only enhances the reality and immersion via physically-based dynamics, but also provides fine-grained interaction and manipulation controllablity.
Although all the study participants appreciated the efficiency and effectiveness of our proposed VR-GS, it still remains to be improved. Firstly, rendering high-fidelity Gaussian kernels in VR is computationally demanding. As a result, rendering generative dynamics in a large scene with 2K resolution might lead to potential latency issues in our system. Secondly, we will extend our system to fluid simulation as a future work, a significant aspect of simulation and editing in realistic scenarios. In future iterations, we aim to incorporate a broader range of materials, such as fluid and cloth, to enhance the system’s capabilities.

### v2

@@ 7. User Study
We conducted a user study including two tasks with 10 participants: 2 professionals (P1–P2) with 5 and 7 years of experience in 3D VFX software (including Houdini and Blender), and 8 novices (P3–P10), with P2 also having 2 years of VR experience. Our study used the hardware specified in the experimental section and included the Fox, Bear, House, Ring Toss, and Toy Collection demos. Participants received a 10-minute video tutorial on our system before completing two tasks. After task 2, they answered a questionnaire covering usability (System Usability Scale [Bangor et al., 2009]) and subjective feedback on individual and overall system features, as shown in Section 7.2. Additionally, we conducted a 20-minute semi-structured interview for more in-depth feedback.
@@ 7.1. Tasks
The participants were required to play two VR games developed by our system: ring toss and toy collection (Figure 11) in two different settings (S1: physics-based and S2: transform-based interaction). Each game had a 5-minute duration under each configuration. A 5-minute break was allowed between sessions. The arrangement of sessions and tasks used a Latin square design to counter potential learning effects. We requested users to rate the immersive and realistic experience in VR using a 7-point Likert scale.
@@
The participants are asked to edit the scene and generate dynamics freely in aforementioned scenes (shown in Figure 9), Figure 10 and Figure 11) with our prototype system, which includes geometry-based editing, physics-based editing, transform, rotation, duplicating, undoing, rescaling, etc. Task 2 is designed to evaluate the detailed potential factors impacting VR immersive experience.
@@ 7.2. Results and Discussion
According to a paired t-test conducted on the score collected from task 1, it could be noticed that physics-based interaction significantly enhanced user immersion and realism compared to transform-based interaction (physics-based: 6.1 vs transform-based: 4.8 on average, p = .0227). As shown in Figure 14, our system has received overall positive opinions. Users spoke highly of the placement of virtual content. Transforming objects without physics rules resulted in floating objects, while in our system, all virtual content placements, such as putting toys onto a sofa, are followed by physics rules and are consistent with the user’s understanding. P7 commented, “I love this a lot. I have a dog. Petting the fox is really like what I did to my dog at home. I felt so real. Moreover, the placement of the toy is awesome. After they all fell to the basket, the basket even shook for a while.” The VR-GS system’s realistic lighting and physics-based dynamics create an immersive experience. As a professional user, P2 spoke highly of high-fidelity generative dynamics and illumination. “Those motions of virtual content are just like what I see in the physical world. I can’t wait to take photos of my own house and then put them in VR with my video game character.” Users rated the system highly on ease of use (4.6/5) and overall satisfaction (4.8/5). With a System Usability Scale (SUS) score of 83.5, VR-GS is classified as “excellent” according to [Bangor et al., 2009].
@@ 8. Conclusion and Future Work
We presented a physical dynamics-aware interactive Gaussian Splatting system for addressing challenges in editing real-time high-fidelity virtual content. By leveraging the advancements in Gaussian Splatting, VR-GS bridges the quality gap traditionally observed between machine-generated and manually created 3D content. Our system not only enhances the realism and immersion via physically-based dynamics, but also provides fine-grained interaction and manipulation controllability.
Although all the study participants appreciated the efficiency and effectiveness of our system, it still remains to be improved. Firstly, rendering high-fidelity Gaussian kernels in VR is computationally demanding. As a result, rendering generative dynamics in a large scene with 2K resolution might lead to potential latency issues in our system. Secondly, the physical parameters in our system are manually defined. Estimating parameters from videos like PAC-NeRF [Li et al., 2023a], or leveraging large-vision models, would be interesting directions to automate this process. In future work, we aim to incorporate a broader range of materials, such as fluid [Feng et al., 2024a] and cloth, to enhance the system’s capabilities. Furthermore, it is also interesting to explore how to utilize large multimodal models to assess the fidelity of the generated dynamics.
Acknowledgements.
We thank the anonymous reviewers for their valuable feedback. We acknowledge support from NSF (2301040, 2008915, 2244651, 2008564, 2153851, 2023780), UC-MRPI, Sony, Amazon, and TRI.